What is the pKb for Pyridine, if the value of Kb is 7.7 x 10-9?


1
Expert's answer
2021-08-23T02:56:32-0400

Solution:

pKb = - log(Kb)

pKb = - log(7.7×10-9) = -(-8.11) = 8.11

pKb = 8.11


Answer: The pKb for pyridine is 8.11